# PRISM-DP: Spatial Pose-based Observations for Diffusion-Policies via Segmentation, Mesh Generation, and Pose Tracking

### v2

@@ 4.3 Real-World Experiments
Figure 6: Success rates (SR) and training time per epoch (TE) for all real-world tasks. The evaluated baselines and variants are the same as simulation experiments except for policies learned from ground-truth states since ground-truth states are no long accessible in real world.
@@
Results. Results for real-world experiments are reported in Fig. 6. Consistent with simulation findings, policies trained using estimated poses from generated meshes perform on par with those using poses from ground-truth meshes, and overall outperform all other baselines in both tasks. In contrast, small-capacity image-based diffusion policies fail completely across both tasks. This likely stems from the increased noise and complexity in real-world RGB observations, which overwhelms their limited representational capacity. Larger image-based diffusion policies achieve marginally lower performance than pose-based policies in the block stacking task, but show a more pronounced performance gap in the drawer interaction task. These results corroborate our simulation findings and reinforce the conclusion that scaling image-based policies to handle complex tasks is increasingly challenging.
5 Discussion
@@
In this work, we introduce PRISM-DP, a novel diffusion policy framework that integrates segmentation, mesh generation, pose estimation, and pose tracking models to enable efficient policy learning from estimated object poses. By leveraging structured pose-based observations instead of high-dimensional RGB inputs, PRISM-DP significantly improves both task performance and training efficiency. Our experiments demonstrate that PRISM-DP not only outperforms image-based diffusion policies at comparable model scales but also surpasses much larger image-conditioned baselines. These results highlight the importance of structured, low-dimensional representations for scaling up diffusion-based visuomotor policies in robotics.
To our knowledge, this is the first work to demonstrate that diffusion policies conditioned on pose-based observations can consistently and significantly outperform those conditioned on raw image observations in real-world settings on suitable tasks. While our current approach has several limitations that may hinder its immediate broad deployment (outlined below), this initial evidence alone underscores the potential of structured, pose-based observations as a promising direction for future research on diffusion policies or related imitation learning paradigms.
Limitations. Our work has several limitations that suggest directions for future research. First, current mesh generation methods are either too computationally expensive or fail to produce meshes suitable for reliable pose estimation. The Meshy model [42] used in this work, for instance, requires approximately one minute to generate a single mesh. Alternative methods we tested [59, 38, 35] were unable to produce meshes of sufficient fidelity for downstream pose estimation. Although generated meshes often achieve photorealistic appearance, their topology deviates substantially from ground-truth quality. For example, the generated can shown in Fig. 5 contains approximately $1.5\times 10^{5}$ triangular faces, whereas the corresponding ground-truth mesh has only 956. This excessive triangle count inflates memory usage and increases rendering time during pose estimation and tracking, resulting in latency bottlenecks. Improving the speed, fidelity, and topological quality of mesh generation remains an important direction for future work. Promising approaches may include model distillation, mesh decimation, the use of geometric inductive biases, and incorporation of explicit topological constraints to produce compact yet accurate 3D models that better support efficient pose estimation.
Our current system also assumes that once a target object enters the camera’s field of view, it remains visible throughout task execution. This assumption was not overly restricting in our relatively uncluttered experimental environments but would likely break down in more complex, dynamic settings where occlusions are common. Moreover, our current method for initializing object poses relies on a human-in-the-loop procedure: when a target object first appears, the user must manually pause execution and provide point prompts to SAM2 through a GUI to generate an initial segmentation mask for pose estimation. While effective, this procedure limits scalability to environments with many objects or frequent changes, and reducing this dependence on human input remains an important challenge.
Additionally, our approach is restricted to handling rigid-body motion. It cannot model nonrigid object deformations, such as folding a shirt or manipulating flexible materials–scenarios where, in theory, image-based observation may still succeed at recognizing necessary patterns.
Another limitation stems from our use of unit quaternions to represent the orientation of target objects. While quaternions provide a broadly applicable starting point for reporting on initial results, prior work suggests that alternative rotation representations, such as continuous 6D parameterizations [60], can offer improved learning stability and performance. Exploring more structured orientation parameterizations is an important direction for future improvements.
Overall, while our method demonstrates strong and promising initial performance, addressing these limitations will be critical for scaling to more complex, cluttered, and dynamic real-world environments.
Appendix A Appendix
A.1 Diffusion Policy Training Details
Conditioned on $\mathbf{o}_{t-H_{o}:t}$ that includes the estimated poses and robot states, the policy $\pi_{\theta}$ is implemented as a diffusion policy that outputs a predicted sequence of actions $\mathbf{a}_{t:t+H_{p}}$. For a diffusion model in general [33], due to Markov property, the forward noising process can be reparameterized using a closed-form solution that directly samples $\mathbf{x}_{k}$ from $\mathbf{x}_{0}$
where $\boldsymbol{\varepsilon}$ is a standard Gaussian noise vector sampled with mean 0 and variance 1, $\bar{\alpha}_{k}$ is a variance scheduler parameter of the noise scheduler, and $k$ is the timestep for denoising or introduce noise.
Intuitively, $\bar{\alpha}_{k}$ controls how much of the original clean signal $\mathbf{x}_{0}$ is preserved versus how much noise is introduced at timestep $k$. As more steps of noise are added by increasing $k$, $\bar{\alpha}_{k}$ generally decreases, causing $\mathbf{x}_{k}$ to become progressively noisier.
During inference, the goal is to reverse this forward diffusion process. Instead of recovering $\mathbf{x}_{0}$ directly in a single step, the model learns to gradually denoise $\mathbf{x}_{k}$ through an iterative procedure. Specifically, at each timestep $k$, the model predicts the noise component $\boldsymbol{\varepsilon}$ present in the current noisy sample $\mathbf{x}_{k}$, using a learned denoising network $\epsilon_{\theta}(\mathbf{x}_{k},k,\mathbf{y})$, where $\mathbf{y}$ represents conditioning information such as observations.
Given this predicted noise, the model constructs the mean $\mu_{\theta}(\mathbf{x}_{k},k,\mathbf{y})$ of the reverse Gaussian distribution $p_{\theta}(\mathbf{x}_{k-1}\mid\mathbf{x}_{k},\mathbf{y})$ used to sample the denoised version $\mathbf{x}_{k-1}$ at the previous timestep:
where $\alpha_{k}$ is another variance-related parameter derived from the noise schedule. Specifically, $\alpha_{k}=1-\beta_{k}$ where $\beta_{k}$ is the variance increment at step $k$.
The model $\epsilon_{\theta}$ is trained to predict $\boldsymbol{\varepsilon}$ given $\mathbf{x_{k}}$, the denoising timestep $k$, and the conditioning $y$ by minimizing the mean squared error loss between the predicted noise by $\epsilon_{\theta}$ and the ground-truth noise $\boldsymbol{\varepsilon}$
$\mathcal{L}(\theta)=\mathbb{E}_{\mathbf{x}_{0},\boldsymbol{\varepsilon},k}\Bigl{\|}\boldsymbol{\varepsilon}-\epsilon_{\theta}(\mathbf{x}_{k},k,\mathbf{y})\Bigr{\|}^{2}$
In PRISM-DP, we instantiate this diffusion model $\epsilon_{\theta}$ as our policy $\pi_{\theta}$ on action sequences, replacing $\mathbf{x}_{k}$ with $\mathbf{a}_{t:t+H_{p},k}$, and conditioning on the observation window $\mathbf{y}=\mathbf{o}_{t-H_{o}:t}$. Since each observation concatenates a robot state and object poses, the training objective becomes:
$\mathcal{L}_{\text{PRISM}}(\theta)=\mathbb{E}_{\mathbf{a}_{t:t+H_{p},0},\boldsymbol{\varepsilon},k}\left\|\boldsymbol{\varepsilon}-\epsilon_{\theta}(\mathbf{a}_{t:t+H_{p},k},k,[\mathbf{s}_{t},\left\{\hat{\mathbf{T}}^{t}_{i}\right\}_{i=1}^{J}])\right\|^{2}.$

### v3

@@ IV-C1 Experimental Settings
Figure 6: Success rates (SR) and training time per epoch (TE) for all real-world tasks. The evaluated baselines and variants are the same as simulation experiments except for policies learned from ground-truth states since ground-truth states are no longer accessible in the real world.
IV-C2 Results
@@
. Results for real-world experiments are reported in Fig. 6. Consistent with simulation findings, policies trained using estimated poses from generated meshes perform on par with those using poses from ground-truth meshes, and overall outperform all other baselines in both tasks. In contrast, small-capacity image-based diffusion policies fail completely across both tasks. This likely stems from the increased noise and complexity in real-world RGB observations, which overwhelms their limited representational capacity. Larger image-based diffusion policies achieve marginally lower performance than pose-based policies in the block stacking task, but show a more pronounced performance gap in the drawer interaction task. These results corroborate our simulation findings and reinforce the conclusion that scaling image-based policies to handle complex tasks is increasingly challenging.
V Discussion
@@
In this work, we introduced PRISM-DP, a framework integrating segmentation, mesh generation, and pose tracking for efficient diffusion policy learning from estimated object poses. By using structured, low-dimensional pose observations instead of high-dimensional RGB images, PRISM-DP significantly improves both task performance and training efficiency. It outperforms image-based policies of a similar scale and surpasses much larger image-conditioned baselines, highlighting the value of structured representations. To our knowledge, this is the first work to show that diffusion policies conditioned on pose-based observations can consistently outperform those conditioned on raw images in real-world settings.
Despite these promising results, our approach has limitations. The generated meshes, while visually adequate, often have a suboptimal topology with excessive triangles, which may increase memory usage and latency during pose tracking. Additionally, our method is restricted to rigid-body motion and cannot model non-rigid deformations like folding cloth. Addressing these challenges will be critical for scaling this approach to more complex and dynamic environments.